OpenStudy (anonymous):

The formula for the perimeter of a rectangle is P = 2l + 2w. Find w if P = 68 and l = 10.

OpenStudy (anonymous):

w=24

OpenStudy (eyust707):

want me to show you why?

OpenStudy (anonymous):

its pretty obvious

OpenStudy (anonymous):

plug an chug

OpenStudy (eyust707):

its obvious that it isnt obvious to isabella

OpenStudy (anonymous):

p = 2l + 2w 68 = 2(10) +2w 68 - 20 = 2w 48 = 2w w=24

OpenStudy (eyust707):

it's easy for you and I.. the goal is to make it that way for her ;)

OpenStudy (anonymous):

You don't have to be so darn rude, some people take time to learn something. Sorry I am one of them. But yes, can you please show me how? (:

OpenStudy (eyust707):

ok so do you get the first step where we just put in the number for the letters?

OpenStudy (anonymous):

i showed u

OpenStudy (eyust707):

talk about obvious.. lol obviously she hasn't learned how to perform all those steps

OpenStudy (eyust707):

so im gunna teach her =)

OpenStudy (anonymous):

yes. i get that part. 68=2*10+2w

OpenStudy (eyust707):

ok then multiply 2*10

OpenStudy (eyust707):

what do you get?

OpenStudy (anonymous):

20

OpenStudy (eyust707):

so whats the equation now look like?

OpenStudy (anonymous):

68 = 20 - 2w

OpenStudy (eyust707):

yep then subtract 20 from both sides